Please solve for green boxes and please show excel formulas.Risk and Return Porfolio Weights and the Security Market Line Do not round any calculations From the Topic Application of th
Answer #1
Dear Student
Thank you for using homeworklib
Please find below the answer
Statement showing Computations
Particulars Amount
Let weight of stock A be w
Let weight of stock B be 1-w
1.1 = 1.3*w + .85 (1-w)
1.1 = 1.3w=.85 - .85w
.25 = .45w
w = .25/.45 = .5556
1-w = 1-.5556 = .4444
Weight of stock A=                      0.5556
Weight of stock B                    0.4444
Portfolio Expected return = 7.80%*.5556 + 5.80%*.4444
Portfolio Expected return = 4.33%+ 2.58% 6.91%
Stock A
7.80% = Rf + 1.3*MRP
Rf = 7.80% - 1.3MRP
Stock B
5.80% = Rf + .85*MRP
Rf = 5.80% - .85MRP
7.80% - 1.3MRP = 5.80% - .85MRP
2.00% =.45MRP
MRP = 2.0%/.45
Market risk premium 4.44%
Rf = 5.80% - .85*4.44% 2.03%
Risk free rate is 2.03%